OpenStudy (anonymous):

Solve for x. lnx(x^2-3x-4)+lnx(x^2-9x+20)=0

OpenStudy (anonymous):

@Mertsj please help!

OpenStudy (anonymous):

couldnt you factor the x out as well?

OpenStudy (anonymous):

whered the e^0 come from?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

wait what? sorry i was just guessing about the base i dont know..

OpenStudy (dumbcow):

after factoring, you could also say this \[\ln x^{2}(x+1)(x-4)^{2}(x-5) = 0\]

OpenStudy (dumbcow):

it doesn't matter what the base is....a^0 = 1

OpenStudy (mertsj):

I will leave you in the capable hands of dumbcow as too many cooks spoil the broth. Good Luck.

OpenStudy (dumbcow):

haha

OpenStudy (anonymous):

lnx(x^2-3x-4)+lnx(x^2-9x+20)=0 whats next after this?

OpenStudy (anonymous):

umm i meant to type what mertjs put but he deleted his answer :/

OpenStudy (dumbcow):

well basically you have to find the roots of a 6th degree polynomial \[x^{2}(x-4)^{2}(x+1)(x-5) = 1\] from analyzing the graph, you know that solution is less than -1 and greater than 5